OpenStudy (anonymous):

(sin^2x)(cos^2x) use power reducing formulas to write the expression in terms of the first power of cosine

OpenStudy (loser66):

\[sin^2x = \frac{1-cos2x}{2}\] \[cos^2x=\frac{1+cos2x}{2}\] ------------------------------- \[sin^2x *cos^2x = timethemtogether\]

OpenStudy (anonymous):

@Loser66 but don't you get cos^2 again?

OpenStudy (loser66):

I know what you mean but why do you expand them to get that form? let (...)*(...) . so? no power there, right?

OpenStudy (anonymous):

@Loser66 so does it become cos4x instead of what I was thinking like... 2cos^2x?

OpenStudy (loser66):

or if you dare not trap your prof like what I did, you can expand them, and then use the power reduce again.

OpenStudy (loser66):

\[(\frac{1+cos2x}{2})*(\frac{1-cos2x}{2})=\frac{1}{2}*(1-cos^22x)\] now, replace again by the trig of \[cos^2(2x) = \frac{1+cos(4x)}{2}\] can you step up from here?

OpenStudy (loser66):

\[\frac{1}{2} * (1 - \frac{1+cos (4x)}{2})\]got what I mean, friend?

OpenStudy (anonymous):

@Loser66 yeah thanks!

OpenStudy (loser66):

good, but didn't get a beautiful answer yet, need some more steps to simplify it. I don't think it's hard to you. when I trapped my prof, he got mad at me but then, we were closer.